A rocket initially at rest steadily gains speed for 2.80s while traveling upwards 20.0m.

what was the rocket's final speed?


1
Expert's answer
2021-10-08T09:22:04-0400

Let's first find the acceleration of the rocket:


"d=\dfrac{1}{2}at^2,""a=\dfrac{2d}{t^2}=\dfrac{2\cdot20\ m}{(2.8\ s)^2}=5.1\ \dfrac{m}{s^2}."

Finally, we can find the rocket's final speed:


"v=v_0+at=0+5.1\ \dfrac{m}{s^2}\cdot2.8\ s=14.28\ \dfrac{m}{s}."